OpenStudy (anonymous):

Find the area of the surface by rotating the graph of f(x) = sqrt(6-x) about x-axis from x = 0 to x = 4

OpenStudy (anonymous):

@mtbender74

OpenStudy (anonymous):

How would we start? I can't graph anything >.<

OpenStudy (anonymous):

ok...i'm going to build off of the link from bob

OpenStudy (anonymous):

did you look at it?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

so we need to first build the pieces we need...

OpenStudy (anonymous):

ok let's begin

OpenStudy (anonymous):

\[S=\int2\pi(\sqrt{6-x})ds\] so we need to calculate ds first

OpenStudy (anonymous):

\[ds=\sqrt{1+(\frac{dy}{dx})^{2}}dx\]

OpenStudy (anonymous):

so dy/dx is the derivative of your function...let's see if we can get ds...

OpenStudy (anonymous):

what have you gotten for ds?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

I got a double integral. I don't think I did something right.

OpenStudy (anonymous):

What are you getting?

OpenStudy (anonymous):

I've been having problems with the site today...sorry.

OpenStudy (anonymous):

were you able to calculate ds?

OpenStudy (anonymous):

I ended up with a double integral, which I doubt is the correct answer

OpenStudy (anonymous):

no...i got a single integral...let me walkthrough my steps and see if you can go from there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y=(6-x)^{1/2}\] \[\frac{dy}{dx}=\frac{-1}{2(6-x)^{1/2}}\] \[ds=\sqrt{1+(\frac{dy}{dx})^{2}}dx=\sqrt{1+(\frac{-1}{2(6-x)^{1/2}})^{2}}dx\]

OpenStudy (anonymous):

\[ds=\sqrt{1+\frac{1}{4(6-x)}}dx=\sqrt{1+\frac{1}{24-4x}}dx\] \[=\sqrt{\frac{24-4x+1}{24-4x}}dx=\sqrt{\frac{25-4x}{24-4x}}dx\] \[=\frac{\sqrt{25-4x}}{2\sqrt{6-x}}dx\]

OpenStudy (anonymous):

all of that make sense?

OpenStudy (anonymous):

Yeah, so far. should i solve for the one at the bottom?

OpenStudy (anonymous):

no...that's just ds...now we put that into the surface area formula, simplify, and integrate...

OpenStudy (anonymous):

\[S=2\pi\int_{0}^{4}yds=2\pi\int_{0}^{4}(6-x)^{1/2}(\frac{\sqrt{25-4x}}{2(6-x)^{1/2}})dx\]

OpenStudy (anonymous):

the (6-x)^(1/2) terms cancel, the 2 in the denominator cancels the 2 on the outside...

OpenStudy (anonymous):

leaving you with \[S=\pi\int_{0}^{4}(25-4x)^{1/2}dx\]

OpenStudy (anonymous):

can you integrate that?

OpenStudy (anonymous):

I think so. one second

OpenStudy (anonymous):

49pi/3

OpenStudy (anonymous):

hmmm...not what i got can you show your steps?

OpenStudy (anonymous):

when you do the integration, you should get \[-\frac{1}{6}\pi(25-4x)^{3/2}\]which gets evaluated from 0 to 4